Let a=(1,2,3), b=(-5,3,-2), c=(2,-4,1) be three points in R3. Find |2b - a + 3c|.


1
Expert's answer
2021-06-02T14:53:09-0400
"\vec a=(1,2,3)"

"\vec b=(-5,3,-2)"

"\vec c=(2,-4,1)"

"2\vec b-\vec a+3\vec c=2(-5,3,-2)-(1,2,3)+3(2,-4,1)"

"=(-10-1+6,6-2-12,-4-3+3)"

"=(-5,-8, -4)"

"|2\vec b-\vec a+3\vec c|=\sqrt{(-5)^2+(-8)^2+(-4)^2}=\sqrt{105}"